An executive invests $28,000, some at 8% and some at 6% annual interest. If he receives an annual return of $2,000, how much is invested at each rate?
Answer by stanbon(75887)

An executive invests $28,000, some at 8% and some at 6% annual interest. If he receives an annual return of $2,000, how much is invested at each rate?
-----------
Equations:
Quantity:: e + s = 28000
Interest::8e +6s = 2000*100
-------------------------------
Modify for elimination::
6e + 6s = 6*28000
8e + 6s = 200000
--------------------------
Subtract and solve for "e"::
2e = 32000
e = $16000 (amt. invested at eight%)
---------------------
Solve for "s":
s = 28000-16000 = 12000 (amt invested at six%
===========================
Cheers,
Stan H.
===================